OpenStudy (anonymous):

Two runners, one averaging 5 miles per hour and the other averaging 4 miles per hour, start at the same place and run along the same trail. The slower runner arrives at the end of the trail a quarter of an hour after the faster runner. How far did each person run?

OpenStudy (anonymous):

Two given rates, two related times, one equal distance. You can set up two \(distance=rate\times time\) equations and set the distances equal to solve.

OpenStudy (anonymous):

I did that. I just was not getting the correct answers when I try to multiply the LCD into each side of the equation.

OpenStudy (anonymous):

hmm, ok, .. can you show me the equations you used?

OpenStudy (anonymous):

\[20(\frac{ x }{ 5 })+20(\frac{ x }{ 4 })=\]

OpenStudy (anonymous):

What's on the other side of the equals sign? Your equation isn't complete. Also, how does it take into consideration the +0.25 hour of time for the slower runner?

OpenStudy (anonymous):

According to the problem the slower running comes in 1 hour after the fast runner. I just figured out the problem but I don't know how I actually came to the final number. I got 5 miles per hour now. Some how I was not completing it somewhere and I was only get 20.

OpenStudy (anonymous):

Now I'm not sure if we're reading the same question. I'm seeing \(r_A=5\) \(r_B=4\) \(d_A=d_B=d\) \(t_B=0.25+t_A\) Find d. \(d=r*t\) \((r_A)(t_A)=(r_B)(0.25+t_A)\) Let \(t_A=t\) \(5t=4(0.25+t)\) Solve for t, plug in to either distance equation to find distance.